# Tomographic redshift dipole: Testing the cosmological principle

### v2

@@ Methods
All data is available starting from data.sdss.org/sas. The SDSS DR16 data used to perform the eBOSS measurements is available at dr17/eboss/lss/catalogs/DR16, and their respective EZmock catalogs at dr17/eboss/lss/EZmocks/v1_0_0/realistic. The SDSS DR12 BOSS data is accessible at dr12/boss/lss, and their respective MultiDark-Patchy mocks at dr12/boss/lss/dr12_multidark_patchy_mocks.
@@
The python code of the estimator and mock generation is hosted at github.com/pdsferreira/tomographic-redshift-dipole.
Acknowledgments
It is a pleasure to thank Nathan Secrest for useful comments and discussions. Funding: PSF thanks FAPES (Brazil) for financial support. VM thanks CNPq (Brazil) and FAPES (Brazil) for partial financial support. Sci-Com cluster: We acknowledge the use of the computational resources provided by the Sci-Com Lab of the Department of Physics at UFES, which was funded by FAPES. Joint CHE / Milliways cluster: We acknowledge the use of the computational resources of the joint CHE / Milliways cluster, supported by a FAPERJ grant E-26/210.130/2023. HOTCAT cluster: We acknowledge the use of the HOTCAT computing infrastructure of the Astronomical Observatory of Trieste of the National Institute for Astrophysics (INAF, Italy). MultiDark-Patchy mocks: The massive production of all MultiDark-Patchy mocks for the BOSS Final Data Release has been performed at the BSC Marenostrum supercomputer, the Hydra cluster at the Instituto de Fısica Teorica UAM/CSIC, and NERSC at the Lawrence Berkeley National Laboratory. We acknowledge support from the Spanish MICINNs Consolider-Ingenio 2010 Programme under grant MultiDark CSD2009-00064, MINECO Centro de Excelencia Severo Ochoa Programme under grant SEV- 2012-0249, and grant AYA2014-60641-C2-1-P. The MultiDark-Patchy mocks was an effort led from the IFT UAM-CSIC by F. Prada’s group (C.-H. Chuang, S. Rodriguez-Torres and C. Scoccola) in collaboration with C. Zhao (Tsinghua U.), F.-S. Kitaura (AIP), A. Klypin (NMSU), G. Yepes (UAM), and the BOSS galaxy clustering working group. EZmocks: We acknowledge the use of the EZmocks galaxy mock catalogues with redshift evolution and systematics for galaxies and quasars of the final SDSS-IV data release [35]. SDSS-III (BOSS) data: Funding for SDSS-III has been provided by the Alfred P. Sloan Foundation, the Participating Institutions, the National Science Foundation, and the U.S. Department of Energy Office of Science. The SDSS-III web site is http://www.sdss3.org/. SDSS-III is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS-III Collaboration including the University of Arizona, the Brazilian Participation Group, Brookhaven National Laboratory, Carnegie Mellon University, University of Florida, the French Participation Group, the German Participation Group, Harvard University, the Instituto de Astrofisica de Canarias, the Michigan State/Notre Dame/JINA Participation Group, Johns Hopkins University, Lawrence Berkeley National Laboratory, Max Planck Institute for Astrophysics, Max Planck Institute for Extraterrestrial Physics, New Mexico State University, New York University, Ohio State University, Pennsylvania State University, University of Portsmouth, Princeton University, the Spanish Participation Group, University of Tokyo, University of Utah, Vanderbilt University, University of Virginia, University of Washington, and Yale University. SDSS-IV (eBOSS) data: Funding for the Sloan Digital Sky Survey IV has been provided by the Alfred P. Sloan Foundation, the U.S. Department of Energy Office of Science, and the Participating Institutions. SDSS acknowledges support and resources from the Center for High-Performance Computing at the University of Utah. The SDSS web site is www.sdss4.org. SDSS is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration including the Brazilian Participation Group, the Carnegie Institution for Science, Carnegie Mellon University, Center for Astrophysics | Harvard & Smithsonian (CfA), the Chilean Participation Group, the French Participation Group, Instituto de Astrofísica de Canarias, The Johns Hopkins University, Kavli Institute for the Physics and Mathematics of the Universe (IPMU) / University of Tokyo, the Korean Participation Group, Lawrence Berkeley National Laboratory, Leibniz Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Astrophysik (MPA Garching), Max-Planck-Institut für Extraterrestrische Physik (MPE), National Astronomical Observatories of China, New Mexico State University, New York University, University of Notre Dame, Observatório Nacional / MCTI, The Ohio State University, Pennsylvania State University, Shanghai Astronomical Observatory, United Kingdom Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Oxford, University of Portsmouth, University of Utah, University of Virginia, University of Washington, University of Wisconsin, Vanderbilt University, and Yale University.
@@ Acknowledgments
Competing interests
The authors declare no competing interests.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure Extended Data Fig. 1: Tracer distribution footprints. Number count maps for each tracer, presented in Mollweide projection using the HEALPix format with $N_{\rm side}=64$.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure Extended Data Fig. 2: Bias maps in Mollweide projection. Similar to Fig. 4 for the individual cases.
Figure Extended Data Fig. 3: Pearson correlation matrix between $\Delta_{1}$ measurements of different tracers. The correlations were obtained using the BOSS and eBOSS mocks. Correlations between the BOSS and eBOSS measurements are absent because their mocks were obtained independently.
Figure Extended Data Fig. 4: Probability distributions of de-biased results without boost. The points mark the $1\sigma$ confidence intervals.
Figure Extended Data Fig. 5: Complete pipeline used in this work. The dashed region represents the steps using observed data. Orange lines indicate the steps of the null case tests (without boost). Blue lines indicate the steps of the bias measurement.
Figure Extended Data Fig. 6: Comparison of dipole measurements in Cartesian coordinates with previous estimates. The distributions for eBOSS, BOSS, eBOSS+BOSS analyses correspond to those presented in Fig. 5.
Table Extended Data Table 1: Significance of tension of previous estimations against our results.
Dipole $\boldsymbol{\Delta}_{1}$ Significance of tension with
Measurement $|\boldsymbol{v}|$ (km/s) $l(^{\circ})$ $b(^{\circ})$ eBOSS BOSS eBOSS+BOSS
[…]

### v3

@@ Acknowledgments
All data is available starting from data.sdss.org/sas. The SDSS DR16 data used to perform the eBOSS measurements is available at dr17/eboss/lss/catalogs/DR16, and their respective EZmock catalogs at dr17/eboss/lss/EZmocks/v1_0_0/realistic. The SDSS DR12 BOSS data (CMASSLOWZTOT) is accessible at dr12/boss/lss, and their respective MultiDark-Patchy mocks at dr12/boss/lss/dr12_multidark_patchy_mocks.
@@
The python code of the estimator and mock generation is hosted at
github.com/pdsferreira/tomographic-redshift-dipole.
@@
Appendix A Pipeline
Figure 11 illustrates the comprehensive pipeline developed for this study.
Figure 11: Complete pipeline used in this work. The dashed region represents the steps using observed data. Orange lines indicate the steps of the null case tests (without boost), while blue lines indicate the steps of the bias measurement.
Appendix B Fisher’s method
Fisher’s method is utilized to consolidate results from multiple independent tests that assess the same hypothesis [45]. It aggregates the $p$-values from each test into a single test statistic, $X^{2}_{2k}$, through the equation:
$X^{2}_{2k}=-2\sum_{i=1}^{k}\ln p_{i},$ (B.1)
where $p_{i}$ represents the $p$-value for the $i$-th hypothesis test. Assuming the null hypothesis is valid and the $p_{i}$ values (or their corresponding test statistics) are independent, $X^{2}_{2k}$ follows a chi-squared distribution with $2k$ degrees of freedom, $k$ being the count of tests being combined. This can be used to determine the combined $p$-value of all Cartesian components of the measured dipole vector, which we express in $\sigma$ units.